# Likelihood-based Mitigation of Evaluation Bias in Large Language Models

### v3

@@ 1 Introduction
Large Language Models (LLMs) exhibit robust language comprehension and text generation capabilities Anil et al. (2023); OpenAI (2023). Relying on this ability, recent studies Liu et al. (2023); Kocmi and Federmann (2023); Chiang and Lee (2023) have employed LLMs as evaluators for natural language generation tasks, surpassing the performance of existing automatic evaluation methods such as BLEU Papineni et al. (2002) and ROUGE Lin (2004). To assess the quality of a text, LLMs output a comprehensive evaluation score based on criteria such as fluency and meaning.
@@
LLMs generate a text based on the likelihood estimations derived from the training process that aims to maximize the likelihood of their large-scale training data. Consequently, it is intuitively possible that the likelihood of a text influences the generation of its evaluation score. However, the likelihood estimations by LLMs may not necessarily align with the quality of the text. For instance, the likelihood calculated by the LLM fluctuates due to superficial differences in sentences, such as word order and sentence structure, even for sentences with identical meaning Kuribayashi et al. (2020). Such fluctuation of likelihood could negatively impact the evaluation score based on meaning criteria.
@@
The extent of likelihood bias may vary with the evaluation criteria used. For instance, likelihood bias is anticipated to be more pronounced in criteria like data coverage, which is less directly related to likelihood. Conversely, the bias is expected to be less significant in criteria like fluency, which is closely related to likelihood. To verify this characteristic of likelihood bias, we adopt two tasks: data-to-text and Grammatical Error Correction (GEC). We use these tasks because, unlike most existing data Freitag et al. (2021); Guan et al. (2021); Kamalloo et al. (2023), the evaluation data for these tasks include multiple criteria such as fluency, grammar, and data coverage.
Our experimental results show that both evaluators based on GPT-3.5 and Llama2-13B Touvron et al. (2023) indeed suffer from likelihood bias. Moreover, our bias reduction method mitigates likelihood bias, and improves evaluation performance in many cases. For reproducibility, we release our code  [^1] https://github.com/stjohn2007/likelihood_bias.
@@ 2 Method
Following a common methodology in LLM-based evaluation Liu et al. (2023); Chiang and Lee (2023), we calculate the LLM’s evaluation score $\text{Score}_{\text{m}}$ based on the models’ response to a prompt. Specifically, we calculate $\text{Score}_{\text{m}}$ as the expected value over candidate scores (e.g. {1, 2, 3, 4, 5}) based on the probability that models output these scores, following the setting of Liu et al. (2023). Our prompt includes a task description and the evaluation criteria, and several few-shot example instances for in-context learning  [^2] The actual prompts and exact equation we use to calculate the $\text{Score}_{\text{m}}$ are provided in Appendix A.. The reason we use in-context learning is that it is known to stabilize the model. This puts us in a position to quantify the strength of likelihood bias.
@@ 3.1 Settings
From a limited set of tasks with available datasets assessed by humans on multiple criteria, we selected two for our experiments: a) data-to-text, converting RDF data into English sentences, and b) GEC. For data-to-text, we use WebNLG+ Castro Ferreira et al. (2020), which contains 2846 instances. $\text{Score}_{\text{h}}$ is provided by human judges, who rated each instance on five criteria (text structure, relevance, fluency, correctness and data coverage). For GEC, we use the TMU-GFM-Dataset Yoshimura et al. (2020), which contains 4221 instances. $\text{Score}_{\text{h}}$ is provided by human judges, who rated each instance on two criteria (grammar and fluency  [^4] All criteria and their definitions are given in Appendix B. The original GEC dataset contains a third criterion, meaning. However, we exclude this criterion because it does not contribute to the overall evaluation (Yoshimura et al., 2020).). We split each dataset into training and test data at a ratio of 4:1.
@@ 3.2 Measuring Likelihood Bias
data coverage .40 .24 .32 * .16 .20 .24 ${}^{{\dagger}}$25
total (micro) .38 .17 ${}^{{\dagger}}$${}^{{\dagger}}$48 .40 .58 * .46
GEC grammar .46 .24 ${}^{{\dagger}}$24 .48 .45 .54 .46
@@ 4 Related Work
Kocmi and Federmann (2023) employ nine GPT models to evaluate translation tasks and verify that GPT-3.5 and larger models possess sufficient capabilities for task evaluation. Chiang and Lee (2023) show that several LLMs’ evaluation results are consistent with those obtained by expert human evaluation in open-ended story generation and adversarial attacks. They also find that LLM-based evaluators are stable over different prompts and sampling algorithms. Zheng et al. (2023) employ GPT-4 to evaluate the conversation ability of LLMs on two benchmarks: MT-bench and Chatbot Arena. Liu et al. (2023) propose G-EVAL, a framework of LLM-based evaluation. They incorporate chain-of-thoughts (CoT) Wei et al. (2023) and introduce a form-filling paradigm and scoring method for stable and fine-grained evaluation. Although LLM-based evaluators have been employed to evaluate several tasks, such as translation Kocmi et al. (2021), summarization Liu et al. (2023), story generation Chiang and Lee (2023), multi-turn conversation Zheng et al. (2023), in this work, we focus on two tasks, data-to-text and GEC, to inspect likelihood bias on multiple criteria.
@@
Zheng et al. (2023) define self-enhancement bias as the tendency of LLM-based evaluators to favor the answers generated by themselves. Their preliminary experiments indicate the existence of this bias. A similar tendency is also reported by Liu et al. (2023). Also, Zheng et al. (2023) introduce verbosity bias, referring to the tendency of LLM-based evaluators to favor longer, more verbose responses. Saito et al. (2023) propose a metric to measure verbosity bias and find both GPT-4 and GPT-3.5 exhibit this bias according to their metric. In contrast to these findings, we introduce likelihood bias, where LLM-based evaluators evaluate high-likelihood sentences and underrate low-likelihood ones compared to human scores. Notably, this is the first work simultaneously proposing methodologies to quantify and mitigate a bias in LLM-based evaluators.
@@ 5 Conclusion
In future work, we aim to investigate the relationship between other biases in LLM-based evaluators, such as self-enhancement bias Zheng et al. (2023) and verbosity bias Zheng et al. (2023); Saito et al. (2023), and likelihood bias. For instance, we may be able to explain self-enhancement bias through the concept of likelihood bias, as the answers generated by LLMs might have high likelihoods calculated by the models themselves. Additionally, we plan to examine the impact of instruction-tuning and model size on likelihood bias.
@@
While we do not foresee any ethical risks caused by our research, LLMs not only exhibit biased likelihood based on surface-level information such as words and sentence structure but also on information like gender, religion, and race Kaneko et al. (2023); Oba et al. (2023); Anantaprayoon et al. (2023). For instance, LLMs might assign a higher likelihood to “She is a nurse” compared to “He is a nurse”. Reducing likelihood bias could potentially address social bias in evaluators. However, it is worth noting that this work does not investigate such aspects, and this remains a task for future research.
@@ Appendix A LLM evaluation method
As is common in LLM-based evaluation Liu et al. (2023); Chiang and Lee (2023), the model is given a prompt $I$, which includes a task description, the evaluation criteria, and an instance $t$, and then predicts score $\text{Score}_{m}$. We also use in-context learning, with the intention of stabilizing the model. Examples are chosen at random when measuring the bias, and are chosen according to the method described in Section 2.2 when mitigating the bias. Finally, we calculate $\text{Score}_{\text{m}}$ as the expected score over scores. We follow the setting of Liu et al. (2023), who have observed that using the expected score, considering the model’s distribution over scores for each instance, rather than always taking the most likely score, leads to a more robust evaluation. Given score candidates $\{1,2,...,n\}$, the probability of each score $Q(i\mid t,F,I;\theta)$, $\text{Score}_{\text{m}}$ is formulated as follows:
@@
Here, we provide two examples of the prompts used for LLM-based evaluators. Our prompts are inspired by the prompts Liu et al. (2023) used.
@@ Appendix B Dataset
We use WebNLG+ Dataset (CC BY-NC-SA 4.0)  Castro Ferreira et al. (2020). Specifically, we collect instances that have human evaluation scores from their dataset. The total number of instances we use is 2846. We use them following their license. There are five criteria in the original dataset:
@@
We use the TMU-GFM-Dataset (CC BY 4.0)  Yoshimura et al. (2020), which contains 4221 instances. We use them following their license. There are three criteria in the original dataset:
@@
Human annotators rate each instance on these criteria using a 5-point scale from 0 to 4. As mentioned in the footnote, we exclude meaning because, according to the original paper Yoshimura et al. (2020), it does not contribute to the overall evaluation.

### v4

@@ 1 Introduction
Large Language Models (LLMs) exhibit robust language comprehension and text generation capabilities anil2023palm; openai2023gpt4. Relying on this ability, recent studies liu-etal-2023-g; kocmi-federmann-2023-large; chiang-lee-2023-large have employed LLMs as evaluators for natural language generation tasks, surpassing the performance of existing automatic evaluation methods such as BLEU papineni-etal-2002-bleu and ROUGE lin-2004-rouge. To assess the quality of a text, LLMs output a comprehensive evaluation score based on criteria such as fluency and meaning.
@@
LLMs generate a text based on the likelihood estimations derived from the training process that aims to maximize the likelihood of their large-scale training data. Consequently, it is intuitively possible that the likelihood of a text influences the generation of its evaluation score. However, the likelihood estimations by LLMs may not necessarily align with the quality of the text. For instance, the likelihood calculated by the LLM fluctuates due to superficial differences in sentences, such as word order and sentence structure, even for sentences with identical meaning kuribayashi-etal-2020-language. Such fluctuation of likelihood could negatively impact the evaluation score based on meaning criteria.
@@
The extent of likelihood bias may vary with the evaluation criteria used. For instance, likelihood bias is anticipated to be more pronounced in criteria like data coverage, which is less directly related to likelihood. Conversely, the bias is expected to be less significant in criteria like fluency, which is closely related to likelihood. To verify this characteristic of likelihood bias, we adopt two tasks: data-to-text and Grammatical Error Correction (GEC). We use these tasks because, unlike most existing data freitag-etal-2021-experts; guan-etal-2021-openmeva; kamalloo-etal-2023-evaluating, the evaluation data for these tasks include multiple criteria such as fluency, grammar, and data coverage.
Our experimental results show that both evaluators based on GPT-3.5 and Llama2-13B touvron2023llama indeed suffer from likelihood bias. Moreover, our bias reduction method mitigates likelihood bias, and improves evaluation performance in many cases. For reproducibility, we release our code  [^1] https://github.com/stjohn2007/likelihood_bias.
@@ 2 Method
Following a common methodology in LLM-based evaluation liu-etal-2023-g; chiang-lee-2023-large, we calculate the LLM’s evaluation score $\text{Score}_{\text{m}}$ based on the models’ response to a prompt. Specifically, we calculate $\text{Score}_{\text{m}}$ as the expected value over candidate scores (e.g. {1, 2, 3, 4, 5}) based on the probability that models output these scores, following the setting of liu-etal-2023-g. Our prompt includes a task description and the evaluation criteria, and several few-shot example instances for in-context learning  [^2] The actual prompts and exact equation we use to calculate the $\text{Score}_{\text{m}}$ are provided in Appendix A.. The reason we use in-context learning is that it is known to stabilize the model. This puts us in a position to quantify the strength of likelihood bias.
@@ 3.1 Settings
From a limited set of tasks with available datasets assessed by humans on multiple criteria, we selected two for our experiments: a) data-to-text, converting RDF data into English sentences, and b) GEC. For data-to-text, we use WebNLG+ castro-ferreira-etal-2020-2020, which contains 2846 instances. $\text{Score}_{\text{h}}$ is provided by human judges, who rated each instance on five criteria (text structure, relevance, fluency, correctness and data coverage). For GEC, we use the TMU-GFM-Dataset yoshimura-etal-2020-reference, which contains 4221 instances. $\text{Score}_{\text{h}}$ is provided by human judges, who rated each instance on two criteria (grammar and fluency  [^4] All criteria and their definitions are given in Appendix B. The original GEC dataset contains a third criterion, meaning. However, we exclude this criterion because it does not contribute to the overall evaluation (yoshimura-etal-2020-reference).). We split each dataset into training and test data at a ratio of 4:1.
@@ 3.2 Measuring Likelihood Bias
[…]
@@ 4 Related Work
kocmi-federmann-2023-large employ nine GPT models to evaluate translation tasks and verify that GPT-3.5 and larger models possess sufficient capabilities for task evaluation. chiang-lee-2023-large show that several LLMs’ evaluation results are consistent with those obtained by expert human evaluation in open-ended story generation and adversarial attacks. They also find that LLM-based evaluators are stable over different prompts and sampling algorithms. NEURIPS2023_91f18a12 employ GPT-4 to evaluate the conversation ability of LLMs on two benchmarks: MT-bench and Chatbot Arena. liu-etal-2023-g propose G-EVAL, a framework of LLM-based evaluation. They incorporate chain-of-thoughts (CoT) wei2023chainofthought and introduce a form-filling paradigm and scoring method for stable and fine-grained evaluation. Although LLM-based evaluators have been employed to evaluate several tasks, such as translation kocmi-etal-2021-ship, summarization liu-etal-2023-g, story generation chiang-lee-2023-large, multi-turn conversation NEURIPS2023_91f18a12, in this work, we focus on two tasks, data-to-text and GEC, to inspect likelihood bias on multiple criteria.
@@
NEURIPS2023_91f18a12 define self-enhancement bias as the tendency of LLM-based evaluators to favor the answers generated by themselves. Their preliminary experiments indicate the existence of this bias. A similar tendency is also reported by liu-etal-2023-g. Also, NEURIPS2023_91f18a12 introduce verbosity bias, referring to the tendency of LLM-based evaluators to favor longer, more verbose responses. saito2023verbosity propose a metric to measure verbosity bias and find both GPT-4 and GPT-3.5 exhibit this bias according to their metric. In contrast to these findings, we introduce likelihood bias, where LLM-based evaluators evaluate high-likelihood sentences and underrate low-likelihood ones compared to human scores. Notably, this is the first work simultaneously proposing methodologies to quantify and mitigate a bias in LLM-based evaluators.
@@ 5 Conclusion
In future work, we aim to investigate the relationship between other biases in LLM-based evaluators, such as self-enhancement bias NEURIPS2023_91f18a12 and verbosity bias NEURIPS2023_91f18a12; saito2023verbosity, and likelihood bias. For instance, we may be able to explain self-enhancement bias through the concept of likelihood bias, as the answers generated by LLMs might have high likelihoods calculated by the models themselves. Additionally, we plan to examine the impact of instruction-tuning and model size on likelihood bias.
@@
While we do not foresee any ethical risks caused by our research, LLMs not only exhibit biased likelihood based on surface-level information such as words and sentence structure but also on information like gender, religion, and race kaneko2023impact; oba2023contextual; anantaprayoon2023evaluating. For instance, LLMs might assign a higher likelihood to “She is a nurse” compared to “He is a nurse”. Reducing likelihood bias could potentially address social bias in evaluators. However, it is worth noting that this work does not investigate such aspects, and this remains a task for future research.
@@ Appendix A LLM evaluation method
As is common in LLM-based evaluation liu-etal-2023-g; chiang-lee-2023-large, the model is given a prompt $I$, which includes a task description, the evaluation criteria, and an instance $t$, and then predicts score $\text{Score}_{m}$. We also use in-context learning, with the intention of stabilizing the model. Examples are chosen at random when measuring the bias, and are chosen according to the method described in Section 2.2 when mitigating the bias. Finally, we calculate $\text{Score}_{\text{m}}$ as the expected score over scores. We follow the setting of liu-etal-2023-g, who have observed that using the expected score, considering the model’s distribution over scores for each instance, rather than always taking the most likely score, leads to a more robust evaluation. Given score candidates $\{1,2,...,n\}$, the probability of each score $Q(i\mid t,F,I;\theta)$, $\text{Score}_{\text{m}}$ is formulated as follows:
@@
Here, we provide two examples of the prompts used for LLM-based evaluators. Our prompts are inspired by the prompts liu-etal-2023-g used.
@@ Appendix B Dataset
We use WebNLG+ Dataset (CC BY-NC-SA 4.0)  castro-ferreira-etal-2020-2020. Specifically, we collect instances that have human evaluation scores from their dataset. The total number of instances we use is 2846. We use them following their license. There are five criteria in the original dataset:
@@
We use the TMU-GFM-Dataset (CC BY 4.0)  yoshimura-etal-2020-reference, which contains 4221 instances. We use them following their license. There are three criteria in the original dataset:
@@
Human annotators rate each instance on these criteria using a 5-point scale from 0 to 4. As mentioned in the footnote, we exclude meaning because, according to the original paper yoshimura-etal-2020-reference, it does not contribute to the overall evaluation.